# Secure Video Quality Assessment Resisting Adversarial Attacks

### v1

@@ IV-D Ablation study
Fourthly, existing VQA models typically employ straightforward concatenation to integrate different types of embeddings. However, we believe that such concatenation does not fully harness the potential of these sophisticated embeddings. Therefore, SecureVQA designs a fusion module to fuse these embeddings extracted by different kinds of backbone networks through some trainable parameters. Table VIII displays the performance of adopting the fusion module in SecureVQA on clean videos compared with concatenation. It can be observed that after adopting the fusion manner, the performance on these two datasets has been improved to some extent, demonstrating the fusion manner in SecureVQA can enhance the representation capability of the different types of embeddings.
@@ IV-E Computational complexity of the defense strategy
The strategies in SecureVQA have almost no impact on computational complexity and can even speed up training and inference processes compared with previous works. For instance, many previous studies try to extract intra-frame and inter-frame information through the frame with the original resolution [17, 38, 19]. However, In the intra-frame branch in SecureVQA, we employ spatial grid sampling to reduce the resolution of the input frame. In the inter-frame branch, SecureVQA adopts resize operation to reduce the resolution of the input frame and then extract the motion information. Moreover, guardian map also introduces almost no computational complexity. Therefore, the three defense strategies in SecureVQA do not impose significant computational complexity while maintaining effective defense properties.
Besides, SecureVQA designs two branches to handle intra- and inter-frame information, respectively. However, SecureVQA utilizes a relatively large network SlowFast [30] to extract inter-frame information, which requires offline processing due to its higher computational cost. In Table IX, we present the defense effect and the runtime of intra-branch and SecureVQA (intra- &\&& inter-branch) on a single 4090 card under the black-box setting. The experiments are conducted via attacking on the LIVE-VQC dataset, which contains 1080P resolution videos. It can be observed that the runtime of SecureVQA is acceptable when assessing videos with high-resolution videos of 1080P. Moreover, if real-time conditions are the primary focus, only the use of intra-branch can also get satisfactory defense effect. In the future, we may explore the design of a more lightweight network to handle inter-frame information, enhancing its robustness against adversarial videos while maintaining real-time requirements.
@@ V Conclusions
In this paper, we perform a pioneering exploration into the security of VQA and focus on general adversarial defense principles, aiming to endow existing VQA models with security. Building upon our exploration, we propose a secure VQA framework named SecureVQA, which is seemingly simple but can completely resist the current adversarial attack methods against the NR-VQA models while achieving superior VQA performance. Extensive experiments have been conducted, and the following conclusions can be drawn. Firstly, randomness serves as a formidable opponent to the optimized-based attack algorithms. Therefore, we introduce two principles to increase the randomness of the model for the intra-frame defense, namely spatial grid sampling and piexl-wise randomization. Secondly, temporal information is a strong guarantee against intra-frame adversarial perturbations. Therefore, we extract temporal embeddings for the inter-frame defense. The ablation study demonstrates that the principles in our framework can serve as a general component for leading VQA models. These findings from SecureVQA are expected to yield valuable insights for the establishment of secure VQA paradigm.
It is worth noting that the vulnerability of QA models has attracted more and more attention in the last two years. This paper explored the general defense principles from the intra- and inter-frame perspectives. Our empirical results show that the intra-frame defense performs better than the inter-frame defense. This is because current adversarial attacks only focus on spatial perturbation, and no attempt has been made to attack the temporal domain of motion video so far. To this end, we extract the temporal information as a complementary safeguard. However, it might be invalid to the next generation of VQA where the motion information could be exploited for an adversarial attack in the future. We intend to evolve them into learnable mechanisms, anticipating that the guardian map will possess enhanced defensive capabilities. Besides, our two principles (i.e. spatial grid sampling and pixel-wise randomization) contribute to the security of VQA models the most, but, are also at the cost of a slight degradation in VQA accuracy. How to efficiently exploit these two characteristics needs more detailed and deeper exploration. Looking back on widespread social effects, we are the first to defend the adversarial attack on VQA models. Targeting our defense principles, a new round of adversarial attacks could be triggered. Hence, such a “cat-mouse” game has emerged in the VQA field, resulting in a never-ending battle. Despite this, we believe that SecureVQA paves a new avenue toward establishing a reliable and practical QA system since these general-purpose principles outlined in our framework can be seamlessly integrated into existing VQA models.

### v2

@@ IV-D Ablation study
Fourthly, existing VQA models typically employ straightforward concatenation to integrate different types of embeddings. However, we believe that such concatenation does not fully harness the potential of these sophisticated embeddings. Therefore, SecureVQA designs a fusion module to fuse these embeddings extracted by different kinds of backbone networks through some trainable parameters. Table VIII displays the performance of adopting the fusion module in SecureVQA on clean videos compared with concatenation. It can be observed that after adopting the fusion manner, the performance on these two datasets has been improved to some extent, demonstrating that the fusion manner in SecureVQA can enhance the representation capability of the different types of embeddings.
@@ IV-E Computational complexity of the defense strategy
The strategies in SecureVQA have almost no impact on computational complexity and can even speed up training and inference processes compared with previous works. For instance, many previous studies try to extract intra-frame and inter-frame information through the frame with the original resolution [21, 42, 24]. However, In the intra-frame branch in SecureVQA, we employ spatial grid sampling to reduce the resolution of the input frame. In the inter-frame branch, SecureVQA adopts a resize operation to reduce the resolution of the input frame and then extracts the motion information. Moreover, the guardian map also introduces almost no computational complexity. Therefore, the three defense strategies in SecureVQA do not impose significant computational complexity while maintaining effective defense properties.
Besides, SecureVQA designs two branches to handle intra- and inter-frame information, respectively. However, SecureVQA utilizes a relatively large network SlowFast [35] to extract inter-frame information, which requires offline processing due to its higher computational cost. In Table IX, we present the defense effect and the runtime of intra-branch and SecureVQA (intra- &\&& inter-branch) on a single 4090 card under the black-box setting. The experiments are conducted via attacking on the LIVE-VQC dataset, which contains 1080P resolution videos. It can be observed that the runtime of SecureVQA is acceptable when assessing videos with high-resolution videos of 1080P. Moreover, if real-time conditions are the primary focus, only the use of intra-branch can also get a satisfactory defense effect. In the future, we may explore the design of a more lightweight network to handle inter-frame information, enhancing its robustness against adversarial videos while maintaining real-time requirements.
@@ V Conclusions
In this paper, we perform a pioneering exploration into the security of VQA and focus on general adversarial defense principles, aiming to endow existing VQA models with security. Building upon our exploration, we propose a secure VQA framework named SecureVQA, which is seemingly simple but can completely resist the current adversarial attack methods against the NR-VQA models while achieving superior VQA performance. Extensive experiments have been conducted, and the following conclusions can be drawn. Firstly, randomness serves as a formidable opponent to the optimization-based attack algorithms. Therefore, we introduce two principles to increase the security of the model for the intra-frame defense, namely spatial grid sampling and pixel-wise randomization. Secondly, temporal information is a strong guarantee against intra-frame adversarial perturbations. Therefore, we extract temporal embeddings for the inter-frame defense. The ablation study demonstrates that the principles in our framework can serve as a general defense component for leading VQA models. These findings from SecureVQA are expected to yield valuable insights for the establishment of a secure VQA paradigm.
It is worth noting that the vulnerability of QA models has attracted more and more attention in the last two years. This paper explored the general defense principles from the intra- and inter-frame perspectives. Our empirical results show that the intra-frame defense performs better than the inter-frame defense. This is because current adversarial attacks only focus on spatial perturbation, and no attempt has been made to attack the temporal domain of motion video so far. To this end, we extract the temporal information as a complementary safeguard. However, it might be invalid to the next generation of VQA where the motion information could be exploited for an adversarial attack in the future. As we predicted, applying temporal smoothing or filtering, similar to the pixel-wise randomization, might mitigate such attacks. Besides, our two principles (i.e. spatial grid sampling and pixel-wise randomization) contribute to the security of VQA models the most, but, are also at the cost of a slight degradation in VQA accuracy. How to efficiently exploit these two characteristics needs more detailed and deeper exploration. Looking back on widespread social effects, we are the first to defend the adversarial attack on VQA models. Targeting our defense principles, a new round of adversarial attacks could be triggered. Hence, such a “cat-mouse” game has emerged in the VQA field, resulting in a never-ending battle. Despite this, we believe that SecureVQA paves a new avenue toward establishing a reliable and practical QA system since these general defense principles outlined in our framework can be seamlessly integrated into existing VQA models.